OpenStudy (anonymous):

PLEASE IM DESPERATE (6-4)^2=2*6*4+4^2 I need to simplify the function

OpenStudy (godlovesme):

I'll help u :) (6-4)^2= 2* 6* 4 + 4^2 Is that ur question?

OpenStudy (godlovesme):

@nicjohns

OpenStudy (anonymous):

Yes!! Thank you

OpenStudy (anonymous):

@godlovesme

OpenStudy (godlovesme):

My pleasure!! :D Do u have options?

OpenStudy (anonymous):

There aren't any! It's a essay response

OpenStudy (godlovesme):

I worked it out and the statement is false :/ here 2^2= 2*6*4+16 4 isn't equal to 64

OpenStudy (texaschic101):

(6 - 4)^2 = 2 * 6 * 4 + 4^2 2^2 = 2 * 6 * 4 + 16 4 = 2 * 6 * 4 + 16 4 = 48 + 16 4 = 64......um, no....not true

OpenStudy (godlovesme):

Ikr @texaschic101 >.<

OpenStudy (anonymous):

I got the same thing and thought I was doing it wrong

OpenStudy (godlovesme):

Maybe its no solution idk =0 this is complicated y do they give this kinda problems smh

OpenStudy (godlovesme):

Thanks for the medal! :)

OpenStudy (anonymous):

I am doing a collab and she gave me this, an equation she made up and gave me X and Y and told me to simplify the function. It came out to 4=64 and I thought I was incorrect

OpenStudy (anonymous):

No problem!!!!